OpenStudy (anonymous):

are medians of . If AG = (x + 5) in. and GD = (x – 1) in., what is the length of ? A. 6 in. B. 10 in. C. 12 in. D. 18 in.

OpenStudy (anonymous):

*kisses u*

OpenStudy (anonymous):

what

OpenStudy (anonymous):

cant help it ur too beautiful

OpenStudy (anonymous):

uuhh thank you

OpenStudy (perl):

need more info

OpenStudy (anonymous):

it a triangle 

OpenStudy (perl):

your question is incomplete though

OpenStudy (perl):

maybe you could take a picture of the problem and upload it

OpenStudy (anonymous):

AD , BE, AND CF are median of ABC

OpenStudy (perl):

ok so AG = 2/3 * AD

OpenStudy (perl):

x+5 = 2/3 * ( x + 5 + x - 1 )

OpenStudy (anonymous):

it is asking for th length of gd

OpenStudy (perl):

since AD = AG + GD AG = 2/3 * (AG + GD )

OpenStudy (perl):

now substitute

OpenStudy (perl):

I am using the centroid theorem , (the centroid is 2/3 the length from the vertex to the midpoint of the opposite side)

OpenStudy (anonymous):

so it is 18?

OpenStudy (perl):

you need to first find x

OpenStudy (perl):

the answer is GD = 6, actually

OpenStudy (perl):

since AD = AG + GD (segment addition postulate) and because AG = 2/3 * (AD) centroid theorem so putting it together AG = 2/3 * ( AG + GD) AG = 2/3 * AG + 2/3*GD AG - 2/3*AG = 2/3 * GD 1/3*AG = 2/3*GD multiply both sides by 3 AG = 2*GD now substitute x+5 = 2 (x-1)

OpenStudy (anonymous):

thank you

OpenStudy (perl):

so x + 5 = 2x - 2 7 = x now GD = x -1 , so GD = 7-1 = 6